# Probing cluster magnetism with embedded and background radio sources in Planck clusters

### v1

@@ Appendix B Alternative visualisations
In Fig. 18 we show a zoomed-in version of Fig. 2 beyond $a=3R_{\mathrm{500}}$, used to determine the scatter between sources far outside clusters. Additionally, to better visualise the low RM sources in our sample, we plot in Figures 19 and 20 the cluster-induced rotation measure on a logarithmic scale as a function of projected radius and column density, respectively. The electron density profiles of the clusters in our sample, as determined in Andrade-Santos et al. (2017), are shown in Figure 21.
Finally, the $q$ values of the full forward model for $\eta=0.0$ are shown in 23. The best-fit profiles for these fits are shown in Fig 24.
@@ Appendix C Updated polarised source catalogue
Table 3: First 30 rows of the catalogue of 819 polarised radio sources that were detected in this work. The full table is available in electronic form at TODO
@@
 [^6] The columns are identical to Osinga et al. (2022), except the newly added column RRM, which is the observed RM corrected for the Galactic value. If uncertainties are not available, they are not given.

### v2

@@ Appendix B Alternative visualisations
In Fig. 18 we show a zoomed-in version of Fig. 2 beyond $a=3R_{\mathrm{500}}$, used to determine the scatter between sources far outside clusters. We find no significant correlation between RRM and projected radius for sources far outside the central regions, although statistically significant peaks in RRM have been found at large radii in a larger sample of RRMs around cosmic overdensities (Anderson et al. 2024).
To better visualise the low RM sources in our sample, we plot in Figures 19 and 20 the cluster-induced rotation measure on a logarithmic scale as a function of projected radius and column density, respectively. The electron density profiles of the clusters in our sample, as determined in Andrade-Santos et al. (2017), are shown in Figure 21. Finally, the $q$ values of the full forward model for $\eta=0.0$ are shown in 23. The best-fit profiles for these fits are shown in Fig 24.
@@ Appendix C Updated polarised source catalogue
Table 3: First 30 rows of the catalogue of 819 polarised radio sources that were detected in this work.
@@
 [^9] The columns are identical to Osinga et al. (2022), except the newly added column RRM, which is the observed RM corrected for the Galactic value. If uncertainties are not available, they are not given. The full table is available in electronic form at the CDS.